# Unitarity of 4D Lattice Theory of Gravity

### v1

@@ II.1 Euclidean signature
×$\times\mathop{\rm tr}\nolimits\gamma^{5}\bigg{\{}\Omega_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}\Omega_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})}\Omega_{\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})5})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}\bigg{\}}.$ (4)
@@
×$\times\mathop{\rm tr}\nolimits\gamma^{5}\bigg{\{}\hat{\Theta}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})2})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}\bigg{\}},$ (5)
@@
×$\times e^{a}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}e^{b}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})2})}e^{c}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}e^{d}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}.$ (8)
@@
×$\times\prod_{{\cal V}}\int\mathop{\rm d}\nolimits\Psi^{{\dagger}}_{{\cal V}}\mathop{\rm d}\nolimits\Psi_{{\cal V}}\mathop{\rm exp}\nolimits(\mathop{{\mathfrak{A}}}\nolimits).$ (9)
@@
Here the superscript "$t"$ denotes the transposition of the Dirac matrices and spinors. We have:
@@ III.2 Gauge fixing
This number consists of $3(r-2)$ 1-simplices on $\partial K^{3}_{p}=S^{2}$ (31), plus $2r$ simplices $(a_{\alpha}a_{p})$, $(a_{\alpha}a_{p^{\prime}})$, $\alpha=1,\ldots,r$, plus the simplex $(a_{p}a_{p^{\prime}})$. The number of 1-simplices on $\partial K^{3}_{p}$ is calculated as follows. Let $\tilde{n}_{0}=r$, $\tilde{n}_{1}$ and $\tilde{n}_{2}$ be the numbers of 0-,1- and 2-simplices on $S^{2}$, respectively. The Euler characteristic for $S^{2}$ gives the relation $\chi=2=\tilde{n}_{0}-\tilde{n}_{1}+\tilde{n}_{2}=r-\tilde{n}_{1}+\tilde{n}_{2}$. On the other hand, on $S^{2}$ there is the relation $2\tilde{n}_{1}=3\tilde{n}_{2}$. Eliminating $\tilde{n}_{2}$ from the last two equalities, we arrive at the equality of 1-simplices on $S^{2}$: $\tilde{n}_{1}=3(r-2)$. Using equalities (34), (35) and (36) we find the number of cycles on the graph of complex $K^{4}_{pp^{\prime}}$:
@@
×$\times\mathop{\rm exp}\nolimits(i\mathop{{\mathfrak{A}}}\nolimits_{pp^{\prime}})\Phi_{\Sigma}\equiv\mathop{\rm exp}\nolimits(i\mathop{{\mathfrak{A}}}\nolimits_{pp^{\prime}})|\Phi_{\Sigma}\rangle.$ (39)
@@ IV.2.1 Second example
V Conclusion
@@ Acknowledgements.
We are grateful to one of our referees who insisted on the need to study the unitarity problem directly in lattice theory. We are grateful to A.Ya. Maltsev for significant assistance in solving mathematical issues, especially in the field of combinatorial topology. This work was carried out as a part of the State Program 0033-2019-0005.

### v2

@@ II.1 Euclidean signature
$\times\mathop{\rm tr}\nolimits\gamma^{5}\bigg\{\Omega_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}\Omega_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})}\Omega_{\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})5})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}\bigg\}.$ (4)
@@
$\times\mathop{\rm tr}\nolimits\gamma^{5}\bigg\{\hat{\Theta}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})2})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}\bigg\},$ (5)
@@
$\times e^{a}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}e^{b}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})2})}e^{c}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}e^{d}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}.$ (8)
@@
$\times\prod_{{\cal V}}\int\mathop{\rm d}\nolimits\Psi^{{\dagger}}_{{\cal V}}\mathop{\rm d}\nolimits\Psi_{{\cal V}}\mathop{\rm exp}\nolimits(\mathop{{\mathfrak{A}}}\nolimits).$ (9)
@@
Here the superscript $t"$ denotes the transposition of the Dirac matrices and spinors. We have:
@@ III.2 Gauge fixing
This number consists of 3(r2)3(r-2) 1-simplices on $\partial K^{3}_{p}=S^{2}$ (31), plus $2r$ simplices $(a_{\alpha}a_{p})$, $(a_{\alpha}a_{p^{\prime}})$, $\alpha=1,\ldots,r$, plus the simplex $(a_{p}a_{p^{\prime}})$. The number of 1-simplices on $\partial K^{3}_{p}$ is calculated as follows. Let $\tilde{n}_{0}=r$, $\tilde{n}_{1}$ and $\tilde{n}_{2}$ be the numbers of 0-,1- and 2-simplices on $S^{2}$, respectively. The Euler characteristic for $S^{2}$ gives the relation $\chi=2=\tilde{n}_{0}-\tilde{n}_{1}+\tilde{n}_{2}=r-\tilde{n}_{1}+\tilde{n}_{2}$. On the other hand, on $S^{2}$ there is the relation $2\tilde{n}_{1}=3\tilde{n}_{2}$. Eliminating $\tilde{n}_{2}$ from the last two equalities, we arrive at the equality of 1-simplices on $S^{2}$: $\tilde{n}_{1}=3(r-2)$. Using equalities (34), (35) and (36) we find the number of cycles on the graph of complex $K^{4}_{pp^{\prime}}$:
@@
$\times\mathop{\rm exp}\nolimits(i\mathop{{\mathfrak{A}}}\nolimits_{pp^{\prime}})\Phi_{\Sigma}\equiv\mathop{\rm exp}\nolimits(i\mathop{{\mathfrak{A}}}\nolimits_{pp^{\prime}})|\Phi_{\Sigma}\rangle.$ (39)
@@ IV.2.1 Second example
V Discussion
1. Let us ask ourselves the first question that arises when studying the problem of unitarity: will the theory be unitary when the number of degrees of freedom changes in the process of evolution? Let us try to present some (sketchy and shaky) arguments in favor of the existing unitarity of such evolution. Here the argumentation will be fundamentally different from the one given above.
By definition, we will consider a linear operator $\cal U$}, acting in a complex linear space with a scalar product, to be unitary if, as a result of its action on any vector, the scalar product of the resulting vector does not change:
$|B\rangle={\cal U}|A\rangle,\quad\langle B|B\rangle=\langle A|A\rangle.$ (50)
Here we admit that the vector $|B\rangle$ belongs to a space with a changed dimension. This is an expansive interpretation of the concept of unitarity, but it is precisely this interpretation that is needed here.
In our case, we denote the set of bosonic and fermionic variables on 3D complexes $\Sigma$ and $\Sigma^{\prime}$ as $V$ and $V^{\prime}$, respectively, and the transfer matrix transforming the WF $\Phi_{A}\{V\}$ on $\Sigma$ into WF $\Phi_{B}\{V^{\prime}\}$ on $\Sigma^{\prime}$, as $\mathop{\rm exp}\nolimits i\mathop{{\mathfrak{A}}}\nolimits\{V^{\prime},V\}$. The transfer matrix is defined on one 4D lattice layer. Then
$\Phi_{B}\{V^{\prime}\}=\int_{V}\mathop{\rm exp}\nolimits\big(i\mathop{{\mathfrak{A}}}\nolimits\{V^{\prime},V\}\big)\Phi_{A}\{V\},$ (51)
and
$\langle B|B\rangle=\int_{V_{1}}\int_{V_{2}}\int_{V^{\prime}}\mathop{\rm exp}\nolimits\Big(i[\mathop{{\mathfrak{A}}}\nolimits\{V^{\prime},V_{1}\}-\mathop{{\mathfrak{A}}}\nolimits\{V^{\prime},V_{2}\}]\Big)$
$\times\Phi^{{\dagger}}_{A}\{V_{2}\}\Phi_{A}\{V_{1}\}.$ (52)
Since the lattice actions are Hermitian, the inner integral over the variables $V^{\prime}$ is an alternating integral for $V_{1}\neq V_{2}$. The simplest model of an alternating integral is the integral
$q=\left|\int_{0}^{a}e^{i\lambda x}\mathop{\rm d}\nolimits x\right|=\left|\frac{2}{\lambda}\sin\frac{\lambda a}{2}\right|<a.$ (53)
Only for $\lambda=0$ we have $q/a=1$. From this we can conclude that the inner integral (52) saturates at $V_{1}=V_{2}$. In other words, as a result of integration over variables $V^{\prime}$, as the number of these integrations tends to infinity, the integral tends to the integral
$\langle B|B\rangle=\int_{V_{1}}\int_{V_{2}}\Phi^{{\dagger}}_{A}\{V_{1}\}\Phi_{A}\{V_{1}\}{\cal F}\{V_{1}\}\delta_{V_{1}V_{2}},$
$\delta_{V_{1}V_{2}}=0\quad\mbox{for}\quad V_{1}\neq V_{2},\quad\int_{V_{2}}\delta_{V_{1}V_{2}}=1.$ (54)
Indeed, let us denote by $q_{{\cal W}}$ the integral over the variables $V^{\prime}$ of one 4-simplex $s^{4}_{{\cal W}}$ at $V_{1}\neq V_{2}$. According to (53) we have $q_{{\cal W}}/(\mbox{measure volume})<1$. Therefore
$\prod_{{\cal W}}\left(\frac{q_{{\cal W}}}{\mbox{measure volume}_{{\cal W}}}\right)\longrightarrow 0$ (55)
for an increasing number of integrations, that is, for an infinite transfer matrix in space. If we could prove that the functional
${\cal F}\{V_{1}\}=\mathop{\rm const}\nolimits,$ (56)
then the unitarity of the lattice theory would be proven for the case with a non-fixed number of degrees of freedom. Unfortunately, we do not know how to prove Eq. (56) at present. Moreover, it is not clear to us whether this problem is correct.
The fermionic part of the actions in (52) would not cause any significant difficulties in proving unitarity. The problem is in proving the equality (56).
[…]
${\cal L}=e^{\vec{\sigma}\vec{\chi}}=\cosh\chi+\sinh\chi(\vec{\sigma}\vec{n}),$
$\vec{\chi}=\chi\vec{n},\quad\vec{n}=(\sin\theta\cos\phi,\,\sin\theta\sin\phi,\,\cos\phi),$
$0\leq\chi<\infty,$ (57)
and $\vec{\sigma}$ are three Pauli matrices. The gauge invariant measure has the form
$\mathop{\rm d}\nolimits\mu({\cal L})=\frac{1}{12}\mathop{\rm tr}\nolimits\Big\{\big({\cal L}^{-1}\mathop{\rm d}\nolimits{\cal L}\big)\wedge{\cal L}^{-1}\mathop{\rm d}\nolimits{\cal L}\big)\wedge{\cal L}^{-1}\mathop{\rm d}\nolimits{\cal L}\big)\Big\}$
$=\sinh^{2}\chi\sin\theta\mathop{\rm d}\nolimits\chi\wedge\mathop{\rm d}\nolimits\theta\wedge\mathop{\rm d}\nolimits\phi.$ (58)
[…]
$\mathop{\rm d}\nolimits\mu_{\mbox{eff}}=\sinh\chi\mathop{\rm d}\nolimits\chi=\mathop{\rm d}\nolimits\cosh\chi.$ (59)
Since the exponential contains a term of the form $(i\lambda\cosh\chi)$, the integral over variable $\chi$ turns out to be similar to the integral $\int\mathop{\rm d}\nolimits x\mathop{\rm exp}\nolimits(i\lambda x)$. The result of such integration leads to generalized functions that must be integrated during further integration.
3. The question arises about numerical modeling of the theory under study. It seems to us that the Monte Carlo method could be adequate here.
VI Conclusion
@@ Acknowledgements.
We are grateful to one of our referees who insisted on the need to study the unitarity problem directly in lattice theory. We are grateful to A.Ya. Maltsev for significant assistance in solving mathematical issues, especially in the field of combinatorial topology. This work was carried out as a part of the State Program FFWR-2024-0011.